### v3

# SAAS: Solving Ability Amplification Strategy for Enhanced Mathematical Reasoning in Large Language Models

Abstract
Equal Contribution Corresponding
This study presents a novel learning approach designed to enhance both mathematical reasoning and problem-solving abilities of Large Language Models (LLMs). We focus on integrating the Chain-of-Thought (CoT) and the Program-of-Thought (PoT) learning, hypothesizing that prioritizing the learning of mathematical reasoning ability is helpful for the amplification of problem-solving ability. Thus, the initial learning with CoT is essential for solving challenging mathematical problems. To this end, we propose a sequential learning approach, named SAAS (Solving Ability Amplification Strategy), which strategically transitions from CoT learning to PoT learning. Our empirical study, involving an extensive performance comparison using several benchmarks, demonstrates that our SAAS achieves state-of-the-art (SOTA) performance. The results underscore the effectiveness of our sequential learning approach, marking a significant advancement in the field of mathematical reasoning in LLMs.
1 Introduction
The advent of Large Language Models (LLMs) has marked a significant breakthrough in various domains. However, despite their remarkable performance across these domains, a notable challenge persists in the realm of mathematical reasoning Zhao et al. (2023); Lu et al. (2022b); Meadows and Freitas (2022); Qian et al. (2022); Zhou et al. (2022); Lightman et al. (2023); Drori et al. (2021); Zhang et al. (2019). The ability of LLMs to comprehend, interpret, and manipulate mathematical concepts is not yet on par with their linguistic capabilities.
The significance of mathematical reasoning in LLMs involves more than just crunching numbers. It also encompasses the ability to engage in logical thinking, problem-solving, and complex decision-making, which are essential for understanding and generating human-like responses in the different situations Lu et al. (2022b); Meadows and Freitas (2022); Thawani et al. (2021). In other words, mathematical reasoning in LLMs is essential for a comprehensive understanding and manipulation of language in numerous scientific and practical applications. However, the current ability of LLMs in mathematical reasoning hinder their potential in the fields where numerical and logical comprehension are paramount such as coding. Thus, it’s critical challenge to enhance the ability of LLMs in mathematical reasoning.
In this study, we explore a learning approach for enhancing both mathematical reasoning ability and problem-solving ability in LLMs, focusing on learning with both the Chain-of-Thought (CoT) Wei et al. (2022b) and the Program-of-Thought (PoT) Chen et al. (2022); Gao et al. (2023a). The CoT rationale (Figure 1-(a)) consists of a series of intermediate reasoning steps. Although it enhances the reasoning ability of LLMs, it leads to arithmetic calculation errors when dealing with large numbers Chen et al. (2022), resulting a low problem-solving ability. To address this issue, Chen et al. (2022) proposed the PoT rationale (Figure 1-(b)), which expresses the reasoning steps as code and delegate computation steps to an code interpreter. It requires the reasoning steps to be expressed accurately as code. Therefore, we hypothesize that prioritizing the learning of mathematical reasoning ability is helpful for the amplification of problem-solving ability. In other words, the initial learning with CoT is essential for solving challenging mathematical problems, since it improves the mathematical reasoning ability Magister et al. (2022); Shridhar et al. (2023); Jie et al. (2023); Liang et al. (2023).
Our research is motivated by an analysis of existing models Gou et al. (2023); Yue et al. (2023). ToRA Gou et al. (2023) tried to learn reasoning ability as well as PoT by adding reasoning step into the PoT rationale. Similarly, MAmmoTH Yue et al. (2023) tried to learn both CoT and PoT by using both CoT rationale and PoT rationale as training data simultaneously. However, we conjecture that they do not fully utilize the advantages of learning with both CoT and PoT. This is because they did not consider the sequence of CoT learning and PoT learning, resulting a less effective learning.
In this work, we introduce a sequential learning approach, named SAAS (Solving Ability Amplification Strategy), to effectively utilize the strengths of CoT learning and PoT learning. This approach transitions from CoT learning to PoT learning, focusing on enhancing problem-solving ability in PoT learning based on logical skills established in CoT learning. This pedagogical strategy ensures that the competencies developed during CoT learning positively influence the PoT learning phase, leading to an overall improvement in solving challenging mathematical problems.
We validate the rationality and effectiveness of our SAAS via extensive experiments on the reputable benchmarks Cobbe et al. (2021); Hendrycks et al. (2021); Gao et al. (2023b); Patel et al. (2021); Miao et al. (2021); Lu et al. (2022a); Koncel-Kedziorski et al. (2016). Most importantly, SAAS achieved state-of-the-art with remarkable performance. Through this, in this paper, we present a novel and effective perspective (i.e., our hypothesis) within the field of mathematics.
2 Related Work and Background
The field of Large Language Models (LLMs) has witnessed substantial advancements, yet the integration of mathematical reasoning within these models remains a challenging frontier. Existing researches in LLMs primarily focus on the natural language understanding and generation Wei et al. (2022a); Yang et al. (2023), with limited exploration in mathematical problem-solving. The complexity of mathematical problems, which requires not only numerical computation but also logical inference and the understanding of abstract concepts, still remains a notable challenge for LLMs Zhao et al. (2023); Lu et al. (2022b); Meadows and Freitas (2022); Qian et al. (2022); Zhou et al. (2022); Lightman et al. (2023); Drori et al. (2021); Zhang et al. (2019). To address this challenge, many researches are being conducted via the following approaches: 1) prompting approach, 2) fine-tuning approach, and 3) continued pretraining approach.
Prompting Approach
Recent studies are based on the prompting methods for mathematical reasoning without additional training. Recently, the concepts of Chain of Thoughts (CoT) Wei et al. (2022b) and Program of Thoughts (PoT) Chen et al. (2022); Gao et al. (2023a) have emerged as promising approaches to enhance mathematical reasoning in LLMs. The CoT involves breaking down complex reasoning problems into a series of intermediate reasoning steps. This approach has shown promise in improving the accuracy and reliability of LLMs in mathematical problem-solving, by mimicking the human thought process of step-by-step reasoning. However, it is not ideal for solving complex mathematical problems Chen et al. (2022). To address this issue, the PoT introduces a more algorithmic perspective. Specifically, it expresses the reasoning steps as code and delegate computation steps to an code interpreter. This approach allows the LLMs to effectively deal with problems that require a combination of mathematical operations and logical reasoning, by structuring the problem-solving process in a programmatic manner.
Fine-tuning Approach
More recently, many works Luo et al. (2023); Yue et al. (2023); Yu et al. (2023); Gou et al. (2023) focus on the fine-tuning LLMs for mathematical reasoning tasks. WizardMath Luo et al. (2023) proposed Reinforcement Learning from Evol-Instruct Feedback (RLEIF), which integrates supervised fine-tuning (SFT) and proximal policy optimization (PPO) for mathematical reasoning. MAmmoTH Yue et al. (2023) introduces a new hybrid instruction-tuning dataset called MathInstruct [^1] https://huggingface.co/datasets/TIGER-Lab/MathInstruct, which consists of CoT rationale and PoT rationale. MetaMath Yu et al. (2023) proposed a new instruction-tuning dataset named MetaMathQA [^2] https://huggingface.co/datasets/meta-math/MetaMathQA, which is augmented by question bootstrapping methods. ToRA Gou et al. (2023) suggested a series of tool-integrated reasoning agents, which is fine-tuned on the tool-use trajectories (PoT rationale) datasets generated by prompting GPT-4.
Figure 1: Overview of SAAS (Solving Ability Amplification Strategy) with two core strategies: i) sequential learning strategy; ii) cognitive retention strategy.
Continued Pretraining Approach
Some researches Lewkowycz et al. (2022); Azerbayev et al. (2023) continually pretrain a base model to specialize in the mathematical reasoning. Minerva Lewkowycz et al. (2022) is a large language model pretrained on general natural language data and further trained on the scientific and mathematical data. Llemma Azerbayev et al. (2023) was also obtained through continued pretraining Code Llama Roziere et al. (2023) on their own collected data named Proof-Pile-2333https://huggingface.co/datasets/EleutherAI/proof-pile-2.
In this paper, we focus on the fine-tuning approach by integrating the CoT and PoT learning. Motivated by Dong et al. (2023) that showed that the abilities of LLMs can be improved depending on the SFT strategy, we analyze how much performance can be improved depending on the SFT strategy from the perspective of solving challenging mathematical problems.
3 SAAS: Solving Ability Amplification Strategy
In this paper, we hypothesize that learning about the problem-solving ability is more effective after logical skills are well established. To explore this, we propose the sequential learning approach, named SAAS (Solving Ability Amplification Strategy), which transitions from CoT learning to PoT learning as shown in Figure 1. Our SAAS is motivated by the pedagogical strategy of human that first learns logical skills and then develops problem-solving abilities by solving numerous problems Glaser (1984). In the following subsections, we describe CoT learning and PoT learning in details.
Figure 2: Overall procedure of the synthetic data generation.
3.1 Chain-of-Thought Learning
It has been shown in various domains that CoT learning, which trains LLMs with data composed of CoT rationales, improves reasoning ability Jie et al. (2023); Liang et al. (2023). Thus, we first fine-tune the LLM via CoT learning for improving mathematical reasoning ability. The primary objective in this phase is to optimize the model parameters for logically interpreting and responding to mathematical problems.
To achieve this, we employ a widely used optimization approach Yu et al. (2023); Gou et al. (2023) that seeks to find the optimal parameters, denoted as $\theta_{cot}^{*}$, which minimize the negative log-likelihood. This is expressed mathematically as:
$\arg\min_{\theta}-\frac{1}{|D_{cot}|}\sum_{(x_{cot},y_{cot})\in D_{cot}}\log p_{\theta}(y_{cot}|x_{cot}),$ (1)
where $\theta$ represents the learnable parameters of the LLM. The dataset $D_{cot}$ consists of ($x_{cot}$, $y_{cot}$) pairs, where $x_{cot}$ denotes a mathematical question, and $y_{cot}$ is the desired CoT rationale for that question.
This optimization process is designed to ensure that the model learns to generate CoT rationales that are logically consistent throughout the reasoning process. This is particularly important in the field of mathematics, since the rationale behind each step is as critical as the final answer. By minimizing the negative log-likelihood, we effectively guide the model to generate step-by-step explanations that mirror human problem-solving approaches, thus enhancing its overall reasoning capability.
This phase sets the foundation for the subsequent PoT learning phase, where the model’s enhanced reasoning ability, developed through CoT training, is further refined and applied to more complex problem-solving scenarios.
3.2 Program-of-Thought Learning
Although the LLM optimized with parameters $\theta_{cot}^{*}$ demonstrates improved logical skills, it still exhibits limitations in problem-solving ability, particularly in computational accuracy Chen et al. (2022), which will be empirically validated in Section 4.2.4. To amplify this problem-solving ability, building upon the mathematical reasoning established in the CoT learning phase, we further fine-tune the LLM with $\theta_{cot}^{*}$ as its starting point using data composed of PoT rationales.
To accomplish this, we construct a dataset $D_{pot+cot}$ that consists of both PoT and CoT rationales. Notably, we integrate CoT rationales alongside PoT rationales in this dataset. This is because we observed that focusing exclusively on PoT rationales during this phase leads to a deterioration in mathematical reasoning ability in our experiments, as detailed in Table 3. To mitigate this cognitive forgetting, we introduce a cognitive retention strategy. This strategy involves randomly sampling CoT rationales and incorporating them into the PoT learning phase. Such a mixed approach (i.e., congnitive retention strategy) ensures that the LLM retains its previously acquired reasoning skills while adapting to the new learning format.
The objective in this phase is to find the final optimal parameters $\theta^{*}$ of the LLM, which involves minimizing the following negative log-likelihood:
$\arg\min_{\theta_{cot}^{*}}-\frac{1}{|D_{pot+cot}|}\sum_{(x,y)\in D_{pot+cot}}\log p_{\theta_{cot}^{*}}(y|x),$ (2)
where $x$ represents a mathematical question, and $y$ is the desired output, which could be either a PoT rationale or a CoT rationale, for the given question $x$. This approach aims to harmonize the strengths of both CoT and PoT learning, thereby equipping the LLM with enhanced computational accuracy and problem-solving abilities while maintaining its proficiency in logical reasoning.
4 Experiments
In this section, we conduct extensive experiments to answer the following key research questions (RQs):
  • RQ1: Does SAAS quantitatively outperform its competitors for solving challenging mathematical problems?
  • RQ2: Are two core strategies of SAAS (sequential learning, cognitive retention strategy) effective in improving the accuracy?
  • RQ3: Is SAAS effective in solving not only basic but also challenging mathematical problems?
  • RQ4: Does sequential learning that transitions from CoT learning to PoT learning help improve both the mathematical reasoning and computational accuracy?
4.1 Experimental Settings
Seed Dataset Rationale Models Size
MetaMathQA CoT GPT, WizardMath 465K
MATH, GSM8K CoT WizardMath 300K
QANDA CoT WizardMath 120K
MetaMathQA PoT ToRA 60K
MATH, GSM8K PoT ToRA 226K
MathInstruct PoT ToRA 38K
QANDA PoT ToRA 12K
Table 1: Summary of synthetic datasets
Model Size GSM8K MATH GSM-Hard SVAMP TabMWP ASDiv MAWPS Avg.
General Models
GPT-4 - 92.0 45.2 64.7 93.1 67.1 91.3 97.6 78.3
GPT-4 (PAL) - 94.2 51.8 77.6 94.8 95.9 92.6 97.7 86.4
ChatGPT - 80.8 35.5 55.9 83.0 69.1 87.3 94.6 72.3
ChatGPT (PAL) - 78.6 38.7 67.6 77.8 79.9 81.0 89.4 73.3
Claude-2 - 85.2 32.5 - - - - - -
PaLM-2 540B 80.7 34.3 - - - - - -
LLaMa-2 7B 13.3 4.1 7.8 38.0 31.1 50.7 60.9 29.4
Platypus-2 7B 14.4 5.4 8.6 36.7 26.5 47.9 58.4 28.3
CodeLLaMa (PAL) 7B 34.0 16.6 33.6 59.0 47.3 61.4 79.6 47.4
SOLAR-1 10.7B 25.8 8.0 17.1 59.3 33.6 55.1 68.4 38.1
LLaMa-2 13B 24.3 6.3 13.6 43.1 39.5 56.3 70.4 36.2
Platypus-2 13B 23.7 7.1 14.3 50.7 45.3 55.1 69.6 38.0
CodeLLaMa (PAL) 13B 39.9 19.9 39.0 62.4 59.5 65.3 86.0 53.1
CodeLLaMa (PAL) 34B 53.3 23.9 49.4 71.0 63.1 72.4 91.5 60.7
LLaMa-2 70B 57.8 14.4 36.0 73.6 57.5 76.0 92.4 58.2
Platypus-2 70B 45.9 15.0 24.6 74.3 47.3 72.7 91.1 53.0
Mathematics Domain-Specific Models
WizardMath 7B 54.9 10.7 20.6 57.3 38.1 59.1 73.7 44.9
MetaMath 7B 66.5 19.8 - - - - - -
MuggleMATH 7B 68.4 - - - - - - -
Toolformer 7B - - - 29.4 - 40.4 44.0 -
MathCoder 7B 64.2 23.3 - - - - - -
MathCoder-CODE 7B 67.8 30.2 - - - - - -
MAmmoTH 7B 53.6 31.5 - - - - - -
MAmmoTH-CODE 7B 59.4 33.4 - - - - - -
ToRA 7B 68.8 40.1 54.6 68.2 42.4 73.9 88.8 62.4
SAAS 7B 74.3 43.2 58.3 74.3 49.6 77.3 93.6 67.2
ToRA-CODE 7B 72.6 44.6 56.0 70.4 51.6 78.7 91.3 66.5
SAAS-CODE 7B 74.8 45.2 58.1 73.6 64.0 80.4 93.8 70.0
SAAS 10.7B 82.0 50.1 64.9 85.0 72.5 87.5 95.7 76.8
WizardMath 13B 63.9 14.0 28.4 64.3 46.7 65.8 79.7 51.8
MetaMath 13B 72.3 22.4 - - - - - -
MuggleMATH 13B 74.0 - - - - - - -
MathCoder 13B 72.6 29.9 - - - - - -
MathCoder-CODE 13B 74.1 35.9 - - - - - -
MAmmoTH 13B 62.0 34.2 - - - - - -
MAmmoTH-CODE 13B 64.7 36.3 - - - - - -
ToRA 13B 72.7 43.0 57.3 72.9 47.2 77.2 91.3 65.9
SAAS 13B 76.6 46.2 61.6 77.8 58.2 80.5 94.3 70.7
ToRA-CODE 13B 75.8 48.1 60.5 75.7 65.4 81.4 92.5 71.3
SAAS-CODE 13B 79.4 50.6 61.6 80.6 68.2 84.5 95.4 74.3
MathCoder-CODE 34B 81.7 45.2 - - - - - -
MAmmoTH-CODE 34B 72.7 43.6 - - - - - -
ToRA-CODE 34B 80.7 50.8 63.7 80.5 70.5 84.2 93.3 74.8
SAAS-CODE 34B 82.9 52.3 64.1 82.8 73.9 85.4 95.2 76.6
SAAS-Llema 34B 85.4 54.7 67.0 85.2 80.2 87.6 96.6 79.5
WizardMath 70B 81.6 22.7 50.3 80.0 49.8 76.2 86.2 63.8
MetaMath 70B 82.3 26.6 - - - - - -
MuggleMATH 70B 82.3 - - - - - - -
MathCoder 70B 83.9 45.1 - - - - - -
ToRA 70B 84.3 49.7 67.2 82.7 74.0 86.8 93.8 76.9
Table 2: Accuracies of competitors and our SAAS on the mathematical benchmark datasets. Our SAAS models are shown in purple color.
4.1.1 Dataset Details
In this paper, we synthesize GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), MetaMathQA Yu et al. (2023), MathInstruct Yue et al. (2023), and QANDA. The QANDA dataset was gathered manually through direct interaction with the application [^4] https://mathpresso.com/en. The overall procedure of synthetic data generation is illustrated in Figure 2.
Specifically, we synthesize these datasets into Chain-of-Thought (CoT) and Program-of-Thought (PoT) rationales via various models (GPT, WizardMath Luo et al. (2023), ToRA Gou et al. (2023)). To generate diverse synthetic data, we adjust some hyperparameters such as temperature and top_p. Then, we select only the correct responses and eliminate similar ones among these correct responses as in Wang et al. (2022). The detailed descriptions of seed datasets are described in Appendix A. Table 1 provides the summary of our synthetic datasets for fine-tuning.
4.1.2 Training Details
We used the CodeLLaMA 13B model Roziere et al. (2023) as our base model and fine-tuned it with our synthetic datasets by setting the batch size to 128. We set learning rate to $52e-$ 5 and use cosine scheduler with warm-up period (1 epoch). For efficient model training, we used DeepSpeed ZeRO Stage3 Rajbhandari et al. (2020).
4.1.3 Model Details
To evaluate the effectiveness of our SAAS in RQ1, we compared it with several state-of-the-art competitors. They can be divided into the following two groups:
  • General models: GPT-4 Achiam et al. (2023), ChatGPT (gpt-3.5-turbo) OpenAI (2023), Claude-2 Anthropic (2023), PaLM-2 Anil et al. (2023), LLaMA-2 Touvron et al. (2023), Platypus-2 Lee et al. (2023), CodeLLaMA Roziere et al. (2023), SOLAR-1 Kim et al. (2023).
  • Mathematics domain-specific models: WizardMath Luo et al. (2023), MetaMath Yu et al. (2023), MulggleMath Li et al. (2023a), Toolformer Schick et al. (2023), MathCoder Wang et al. (2023), MammoTH Yue et al. (2023), ToRA Gou et al. (2023).
As in Gou et al. (2023), we report CoT prompting results by default, and include PAL Gao et al. (2023a) prompting results for selected models. Within the category of mathematics domain-specific models, WizardMath, MetaMath, and MuggleMath exclusively employ CoT learning for fine-tuning. Conversely, ToRA utilizes solely PoT learning, whereas MathCoder and MammoTh integrate a combination of CoT and PoT learning methodologies for fine-tuning. Also, Toolformer is trained to utilize calculators.
4.1.4 Evaluation Details
We evaluated the model’s performance and its ability to generalize mathematical reasoning using both in-domain and out-of-domain data. For in-domain evaluation, we use the test set of MATH and GSM8K dataset. For out-of-domain evaluation, we utilized the following various datasets, which are used in the previous studies Gou et al. (2023); Yue et al. (2023) and publicly available: GSM-Hard Gao et al. (2023b), SVAMP Patel et al. (2021), ASDIV Miao et al. (2021), TabMWP Lu et al. (2022a), and MAWPS Koncel-Kedziorski et al. (2016) that consists of SingleEQ, SingleOP, AddSub, and MultiArith. These datasets ensure a comprehensive analysis of the model’s applicability across various mathematical contexts.
4.2 Results and Analysis
We highlight the best and the second-best results in each column (i.e., dataset) of the following tables in bold and underline, respectively.
4.2.1 RQ1: Comparison with Competitors
To demonstrate the superiority of our SAAS over competitors, we compare the accuracies of all competitors and SAAS. In this experiment, we utilize LLaMA-2 7B, CodeLLaMA 7B, SOLAR-1 10.7B, LLaMA-2 13B, CodeLLaMA 13B, CodeLLaMA 34B, and Llemma-34B as our base models. [^5] For experiment on the 70B model, we could not proceed it due to hardware constraint.
Table 2 shows the results. We summarize our empirical findings as follows. First, we observed that mathematics domain-specific models outperforms general models with similar size in almost cases. This indicates a requisite for domain-specific models to address complex mathematical problems effectively. Second, among mathematics domain-specific competitors, ToRA, which utilizes solely PoT learning, consistently outperforms all others with similar size, including MathCoder and MammoTH, which integrate a combination of CoT learning and PoT learning methodologies. This implies that simply combining CoT and PoT learning does not effectively solve complex mathematical problems. Therefore, a strategic and careful approach is imperative in the combination of CoT and PoT learning. Third and most importantly, our SAAS consistently and significantly outperforms all competitors with similar size. Specifically, on $\sim$7B size, 7Bsimilar-$\sim$13B size, 13Bsimilar-$\sim$34B size, and 34Bsimilar-$\sim$70B size, SAAS outperforms the best competitors (i.e., ToRA-CODE and ToRA) by up to 5.26%, 7.71%, and 6.28% in terms of average score. Note that although we could not fine-tune 70B model, SAAS with 10.7B showed similar performance to ToRA with 70B. Furthermore, SAAS-Llema demonstrated superior performance than ToRA with 70B. This remarkable performance of SAAS underscore the effectiveness of our sequential learning approach.
Strategy GSM8K MATH
Chain-of-Thought (CoT) 69.7 26.9
Program-of-Thought (PoT) 76.8 47.7
Combination of CoT and PoT 79.0 49.2
SAAS 79.4 50.6
without cognitive retention strategy 79.0 49.6
Reverse SAAS 76.8 47.1
without cognitive retention strategy 69.4 27.6
Table 3: Accuracies of different learning strategies. All improvements are statistically significant with $p$-value $\leq 0.001$.
4.2.2 RQ2: Effectiveness of Sequential Learning and Cognitive Retention Strategy
To further explore what factors contribute to the improvement of our SAAS, we conduct comparative experiments on diverse learning strategies, as shown in Table 3. Specifically, we compare CoT learning, PoT learning, CoT+PoT learning, SAAS that transtions from CoT learning to PoT learning, and reverse SAAS that transtions from PoT learning to CoT learning. In addition, we compare (reverse) SAAS without cognitive retention strategy to validate the effectiveness of this strategy. From Table 3, our empirical findings are summarized as follows:
  1. i)
    Effectiveness of the hybrid learning: Combining of CoT and PoT learning significantly outperforms both CoT learning and PoT learning. This is because CoT learning, which enhances mathematical reasoning ability, and PoT learning, which improves problem-solving ability, play a complementary role;
  2. ii)
    Effectiveness of the sequential learning: Our SAAS without cognitive retention strategy slightly outperforms combining of CoT and PoT learning in MATH only. We conjecture that the absence of significant improvement, despite sequential learning, can be attributed to the deterioration of mathematical reasoning abilities during the PoT learning phase (i.e., cognitive forgetting). Furthermore, reverse SAAS without cognitive retention strategy shows a lower accuracy than combining of CoT and PoT learning. This result indicates that the order of the learning sequences in sequential learning is vital for mathematical reasoning and problem-solving abilities;
  3. iii)
    Effectiveness of the cognitive retention strategy: To mitigate the cognitive forgetting, in Section 3.2, we proposed the cognitive retention strategy, which includes some data samples from first phase in the second phase. (Reverse) SAAS outperforms (reverse) SAAS without cognitive retention strategy, verifying the effectiveness of the cognitive retention strategy.
Figure 3: Accuracies on GSM8K with respect to the number of required reasoning steps.
Figure 4: Responses of different learning approaches for a given question-answer pair.
4.2.3 RQ3: Further Analysis of the Capabilities of SAAS
To analyze the capabilities of SAAS depending on the difficuly of mathemtical problem, we quantitatively assess the break-down accuracies for problems with respect to the reasoning steps as in Shi et al. (2023). Specifically, we segmented the GSM8K dataset into 4 categories based on the number of reasoning steps required to arrive at an answer. Then, we quantified accuracies of CoT learning, PoT learning, and SAAS across each designated category.
As illustrated in Figure 3, SAAS outperforms both CoT learning and PoT learning for all categories. Moreover, as the reasoning steps in a mathematical problem extend (i.e., the difficulty increases), especially the step 5 or above, the difference between our SAAS and other strategies becomes more pronounced. This result supports our hypothesis that prioritizing the learning of mathematical reasoning ability via CoT learning is helpful for the amplification of challenging problem-solving ability.
4.2.4 RQ4: Case Study
To demonstrate that our SAAS is effective in terms of both mathematical reasoning and computational accuracy, we conduct a case study showing the responses of CoT learning, PoT learning, and SAAS for a given question-answer pair. Figure 4 shows the visualization results, where the colored words indicate incorrect responses and the words with no color mark indicate correct responses.
As depicted in Figure 4, CoT learning approach exhibited inaccuracies in arithmetic computations as well as deficiencies in mathematical reasoning. Conversely, PoT approach demonstrated precise calculations yet exhibited a critical deficiency in mathematical reasoning. As we expected, our SAAS exhibited precise computational accuracy along with enhanced mathematical reasoning capabilities (See the more detailed comments than the comments of PoT learning). Through this case study, we demonstrated the following three observations: i) only CoT learning approach leads to arithmetic calculation errors; ii) only PoT learning approach may result in a deficit of mathematical reasoning; iii) sequential learning that transitions from CoT learning to PoT learning help improve computational accuracy as well as mathematical reasoning.
5 Conclusion
In this paper, we demonstrated the following two important points in the sense of solving challenging mathematical problems: (1) prioritizing the learning of mathematical reasoning ability via Chain-of-Thought (CoT) learning is helpful for the amplification of problem-solving ability during Program-of-Thought (PoT) learning; (2) for effective sequential learning, it is necessary to employ a cognitive retention strategy that incorporates some data samples from the initial phase into the subsequent phase. In light of this, we proposed a novel sequential learning approach, named SAAS (Solving Ability Amplification Strategy), which progresses from CoT learning to PoT learning with cognitive retention strategy. Through extensive experiments with the reputable benchmarks, we demonstrated that SAAS consistently and significantly outperforms all competitor, marking a significant advancement in the field of mathematical reasoning in LLMs.
Acknowledgements
This work was supported by the 2023 KT ICT AI2XL Laboratory R&D Fund" project funded by KT.
Limitations
This study, while advancing the field of computational linguistics through the use of Large Language Models (LLMs), encounters several limitations that are important to acknowledge.
Firstly, the intricate nature of LLMs can sometimes lead to unpredictability in their outputs. This unpredictability can be particularly challenging when dealing with mathematical reasoning, where precision and accuracy are paramount, making it difficult to utilize LLMs in applications in the field of mathematics.
Furthermore, despite advancements via our study, LLMs still have limitations in their understanding and application of advanced mathematical concepts. While they can perform well on structured problems, their ability to handle abstract and complex mathematical reasoning is still an area of ongoing research and development.
Additionally, the reliance on synthetic data for training these models also presents a limitation. While synthetic datasets are useful for mitigating the scarcity of real-world data, it may not always accurately capture real-world scenarios, leading to potential gaps in the model’s performance when applied to practical, real-world tasks.
Finally, ethical considerations, particularly around the potential misuse of AI, remain a concern. Ensuring that LLMs are used responsibly and do not perpetuate biases is an ongoing challenge in the field.
In summary, while our study leverages the capabilities of LLMs to enhance mathematical reasoning in computational linguistics, it is important to recognize the limitations related to unpredictability of LLMs, understanding of advanced mathematical concepts, reliance on synthetic data, and ethical considerations. These limitations highlight the need for continued research and development in the field to address these challenges effectively.
Ethics Statement
In this research, we have diligently adhered to the highest ethical standards of scientific inquiry and data management, ensuring the integrity and reliability of our findings. The design and execution of our experiments were grounded in fairness and objectivity, without favoring any particular outcome. This commitment was reflected in our meticulous planning and consistent application of methodologies across various datasets.
We also placed a strong emphasis on data privacy and security, handling all data, especially synthetic data generated for our models, in compliance with relevant data protection laws and guidelines. We confirmed that all the data used in our experiments were free of licensing issues. Our approach to data was characterized by strict anonymization protocols and its use was confined strictly to research purposes. We have strived for transparency in our research process, documenting all methodologies, data sources, and analysis techniques clearly, which underpins our commitment to the reproducibility of scientific research. This allows other researchers to verify our results and build upon our work, contributing to the collective knowledge in the field.
Recognizing the broader impacts of AI and LLMs on society, our research was conducted with a profound sense of responsibility. We were mindful of the ethical implications of AI development and aimed to create models that are effective yet ethically aligned, avoiding any form of biased, discriminatory, or harmful applications of these technologies. We believe our research makes a positive contribution to the field of computational linguistics and AI, particularly in enhancing the mathematical reasoning capabilities of Large Language Models in a manner that is ethically sound and socially responsible.
Our work underscores our commitment to conducting scientifically rigorous and ethically responsible research, maintaining the highest standards of integrity in AI and computational linguistics.
Appendix A Detailed Descriptions of Seed Datasets
The detailed description of each seed dataset is as follows:
  1. i)
    GSM8K Cobbe et al. (2021): It focuses on elementary-level math problems to evaluate abilities that handle logical reasoning and parse and interpret math questions presented in natural language;
  2. ii)
    MATH Hendrycks et al. (2021): It includes a wide range of math problems, ranging from elementary arithmetic to advanced topics such as algebra, calculus, and geometry, which are challenging more than GSM8K;
  3. iii)
    MetaMathQA Yu et al. (2023): It is a dataset augmented through rephrasing question, forward-backward reasoning Jiang et al. (2023), self-verification, and answer augmentation based on GSM8K and MATH;
  4. iv)
    MathInstruct Yue et al. (2023): It consists of a mix of 13 types of CoT and PoT mathematical rationales from various mathematical fields. Specifically, CoT type data consist of GSM8K, GSM8K-RFT Yuan et al. (2023), AQuA-RAT Ling et al. (2017), MATH, THeoremQA Chen et al. (2023) Camel-Math Li et al. (2023b) and College-Math. Otherwise, PoT type data consist of GSM8K, AQuA-RAT, TheoremQA, MathQA Amini et al. (2019) and NumGLUE Mishra et al. (2022);
  5. v)
    QANDA: It consists of a diverse collection of real-world mathematical questions and detailed solutions, catering to a broad spectrum of mathematical concepts and difficulty levels.